# Testing identifying assumptions in Tobit Models

### v2

@@ 2.2. IV Tobit model
Assumption 3(i) states that the instrumental variable is independent of the model’s structural latent variables. The second part of Assumption 3 is the usual instrument relevance in determining the treatment. This assumption also implies that $Z$ is independent of $W,V$, since $W=\alpha_{1}V+U$.
Assumption 4 characterizes the distribution of the latent vector of error terms as bivariate normal. Then, $W,V$ also follows a bivariate normal distribution and $\alpha_{0},\alpha_{1},\gamma_{0},\gamma_{1}$ could be scale normalized so $(W,V)^{\prime}$ follows the standard bivariate normal distribution with covariance $\rho$, i.e., $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. [^7] Note that if $U,V$ follow a bivariate normal distribution, $\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{UV}\\ \rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$, then $W,V$ follow a bivariate normal distribution, $\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}(\mu_{1},\Sigma_{1})$, where $\mu_{1}=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma_{1}=\begin{pmatrix}\alpha_{1}^{2}\sigma^{2}_{V}+\sigma^{2}_{U}+2\rho_{UV}\alpha_{1}\sigma_{V}\sigma_{U}&\alpha_{1}\sigma_{V}^{2}+\rho_{UV}\\ \alpha_{1}\sigma_{V}^{2}+\rho_{UV}&\sigma^{2}_{V}\end{pmatrix}$. Then the reduced form system can be scale normalized. The latter characterization aids the development of the sharp testable implications.
@@ 3. Sharp testable equalities
This section presents the sharp testable equalities for the models described in Section 2. Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, independence, instrument validity, and normality) is valid. The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 4.
@@ 3.1. Classic Tobit model
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq U\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
where $\Phi(.)$ is the standard normal CDF. The first equality follows from the law of total probability. The second through fourth equalities follow from the model’s structure described in Equation 2.4, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y$ for $Y^{*}>0$. The fifth one is given by the latent linear model structure of $Y^{*}$, and finally, the last step follows from assumptions 1 and 2. To cover all possible cases of positive outcomes, note that for any $c_{1}>0$:
$P(c_{1}\leq Y|D=d)$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d).$ (3.2)
Turning to the accumulation point, the observed event of $Y=0$ has probability,
$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi(-\alpha_{0}-\alpha_{1}d).$ (3.3)
The first equality follows from the model’s structure in Equation 2.4, and the second one applies assumptions 1 and 2. The equalities described in equations 3.1-3.3 fully characterize the distribution of $Y$ conditional on $D$, connecting the probabilities in the observed data to those implied by the Tobit model.
Remark 1 (Sharpness).
In the context of model 2.4, equalities 3.1-3.3 are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},D,\tilde{U})$ that satisfies model 2.4, Assumptions 1 and 2, and induces the observed distribution on the data $(Y,D)$. For the proof, see Appendix A.
Then, the equalities 3.1-3.3 are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2. They serve as the basis for the test procedure described in Section 5.
Remark 2 (Non-learnability).
The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable. However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities 3.1-3.3. See Appendix E for the proof.
Remark 3 (Extensions).
The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (Barros et al., 2018), different thresholds (Carson and Sun, 2007), dynamic Tobit models, or including individual specific effects (Wooldridge, 2005; Honore, Kyriazidou, and Powell, 2000; Honoré, 1993). In appendix C, we derive the equalities for the type 2 Tobit model, and similar logic can be applied to other two-part models. Additional testable results for the aforementioned models are discussed in Appendix D.
@@ 3.2. IV Tobit model
We turn our attention to the instrumental variable Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4.
The observed data includes $(Y,D,Z)$, and we characterize the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications. The mapping from observed probabilities to their corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$. Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.4)
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.5)
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$ (3.6)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.7)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
@@
$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.8)
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$

### v3

@@ 2.2. IV Tobit model
Assumption 3(i) states that the instrumental variable is independent of the model’s structural latent variables. The second part of Assumption 3 is the usual instrument relevance in determining the treatment. This assumption also implies that $Z$ is independent of $\tilde{W},\tilde{V}$, since $\tilde{W}=\tilde{\alpha}_{1}\tilde{V}+U$.
[…]
3. Identification
In this section, we provide heuristic arguments for the identification of the parameters of the respective models, expanding on han_identification_2017 and acerenza_testing_2023. These are known in the literature and discussed here for completeness and intuition. The testable implications that are our main result are presented in Section 4.
3.1. Classic Tobit model
Note that,
$1-P(Y=0|D)=1-P(Y^{*}\leq 0|D)=\Phi(\alpha_{0}+\alpha_{1}D).$
The equality follows from the model’s structure in Equation 2.4 and assumptions 1 and 2.
$\Phi^{-1}(1-P(Y=0|D))=\alpha_{0}+\alpha_{1}D.$
The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$. This implies that the ratio of the original parameters is identified since $\frac{\alpha_{1}}{\alpha_{0}}=\frac{\frac{\tilde{\alpha}_{1}}{\sigma}}{\frac{\tilde{\alpha}_{0}}{\sigma}}=\frac{\tilde{\alpha}_{1}}{\tilde{\alpha}_{0}}$. Recall that,
$F_{U|U<c}(u|U<c)=P(U\leq u|U<c)$ $=$ $\frac{P(U\leq\min(u,c))}{P(U<c)}=\begin{cases}\frac{P(U\leq c)}{P(U\leq c)}=1,\text{ for }u\geq c\\ \frac{P(U\leq u)}{P(U\leq c)}=\frac{\Phi(u)}{\Phi(c)},\text{ for }u<c\end{cases}$
Then, $f_{U|U<c}(u|U<c)=\begin{cases}0,\text{ for }u\geq c\\ \frac{\phi(u)}{\Phi(c)},\text{ for }u<c\end{cases}$, and $f_{U|U>c}(u|U>c)=\begin{cases}0,\text{ for }u\leq c\\ \frac{\phi(u)}{1-\Phi(c)},\text{ for }u>c\end{cases}$. Note then that,
$E(Y|D,Y>0)$ $=$ $E(\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U|D,Y>0)$
$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|D,Y>0)$
$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|D,\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U>0)$
$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+E(U|U>-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}D)$
$=$ $\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\frac{\phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}{1-\Phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}$
where the fourth equality uses independence, and the fifth equality uses the properties of standard truncated normal distributions. Since $\alpha_{0},\alpha_{1}$ are identified, then the inverse Mills ratio is identified,
$\frac{\phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}{1-\Phi(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}D)}$ $=$ $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)=\lambda(\alpha_{0}+\alpha_{1}D)$
Thus,
$E(Y|D,Y>0)=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\lambda(\alpha_{0}+\alpha_{1}D)$
Which in turn implies (by the property that we can express $X=E(X|W)+e$, where $E(e|W)=0$):
$Y=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+\sigma\lambda(\alpha_{0}+\alpha_{1}D)+e$
with $e$ mean independent of $D,\lambda(\alpha_{0}+\alpha_{1}D)$. Since $D$ is known and $\lambda(\cdot)$ is identified, we get a three-by-three linear system with a unique solution for $\tilde{\alpha}_{0},\tilde{\alpha}_{1},\sigma$. In other words, the normality assumption allows identification by leveraging the relationship of the variable of interest ($D$) and the outcome at the accumulation point ($Y=0$) to obtain the scaled parameters, and the relationship of treatment with the distribution of $Y$ (when $Y>0$) to overcome the normalization. However, the reliance on the normality and linearity assumptions underlines the importance of adequately testing the model’s structure and assumptions in empirical applications.
3.2. IV Tobit model
Turning the focus to the IV-Tobit model, the first stage is readily identified from the linear structure, $\tilde{\gamma}_{1}=\frac{Cov(Z,D)}{Var(Z)}$, $\tilde{\gamma}_{0}=E(D)-\tilde{\gamma}_{1}E(Z)$. Furthermore, $\tilde{V}=D-E(D)+\frac{Cov(Z,D)}{Var(Z)}E(Z)-\frac{Cov(Z,D)}{Var(Z)}Z$, thus $\tilde{V}$ and $\sigma_{\tilde{V}}^{2}$ are identified. Combining these results implies that we can identify $\gamma_{0}=\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}$ and $\gamma_{1}=\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}$.
Now note that calculating $P(Y=0|Z)$ from the reduce form:
$1-P(Y=0|Z)=\Phi(-\beta_{0}-\beta_{1}Z)$
Where $\beta_{0}=\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}$ and $\beta_{1}=\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}$.
[…]
To identify $\rho$, let $s=c-\beta_{0}-\beta_{1}z$ and $t=d-\gamma_{0}-\gamma_{1}z$. Then, by using
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
$+$ $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\beta_{0}-\beta_{1}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\gamma_{0}-\gamma_{1}z;\rho\right)$
$\equiv$ $\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}},$
We can obtain
$P(Y\leq c,D\leq d|Z=z)$ $=$ $\Phi_{W,V}(s,t;\rho),$
which by classic results of bivariate normal random variables if we differentiate w.r.t. $\rho$ is:
Recall the previous coefficients are normalized, as noted by (wooldridge_econometric_2010, p. 683-84) we can identify the original and reduced form covariance matrices, as well as the true coefficients without the normalization. To explain this in detail, recall that we have identified $\rho$, the covariance between $W,V$ the parameters $\gamma_{1},\gamma_{0},\beta_{1},\beta_{0}$ and the variance of $\tilde{V}$, $\sigma^{2}_{\tilde{V}}$. Similarly to the classic tobit model, we can leverage normality of $W$, independence of $Z$, and exploit the reduced form representation of $Y$ to obtain:
$E(Y|Z,Y>0)$ $=$ $\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\sigma_{\tilde{W}}\lambda(\beta_{0}+\beta_{1}Z)$
Which in turn implies
$Y=\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\sigma_{\tilde{W}}\lambda(\beta_{0}+\beta_{1}Z)+e$
with $e$ mean independent of $(Z,\lambda(\beta_{0}+\beta_{1}Z))$. Since $Z$ is known and $\lambda(\beta_{0}+\beta_{1}z)$ is identified for any given $Z=z$, we can get a solution for $\sigma_{\tilde{W}}$ as one of the coefficients for the linear projection of $Y$ on $(Z,\lambda(\beta_{0}+\beta_{1}Z))$, identifying the non-standardized reduced form error for $Y$. Finally, since $\beta_{0}=\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}=\frac{\tilde{\alpha}_{0}+\tilde{\alpha}_{1}\tilde{\gamma}_{0}}{\sigma_{\tilde{W}}}$, $\beta_{1}=\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}=\frac{\tilde{\alpha}_{1}\tilde{\gamma}_{0}}{\sigma_{\tilde{W}}}$ and we have identified $\sigma_{\tilde{V}},\gamma_{0},\gamma_{1},\tilde{\gamma}_{0},\tilde{\gamma}_{1}$ and $\sigma_{\tilde{W}}$ we can recover $\tilde{\alpha}_{0},\tilde{\alpha}_{1}$.
It now only remains to recover the original variance-covariance structure. Under bivariate normality of the latent variables in the structural model, we can express the following relationships between $U,\tilde{V}$ with $\tilde{W}=\tilde{\alpha}_{1}U+\tilde{V},V$ and $W=\frac{\tilde{\alpha}_{1}U+\tilde{V}}{\sigma_{\tilde{W}}},V=\frac{\tilde{V}}{\sigma_{\tilde{V}}}$
$\begin{pmatrix}U\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}\Bigg(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}\sigma^{2}_{U}&\rho_{U\tilde{V}}\\ \rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}\Bigg)$
$\begin{pmatrix}\tilde{W}\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}\Bigg(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}&\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}\\ \tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}\Bigg)$
$\begin{pmatrix}W\\ V\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\ \end{pmatrix},\begin{pmatrix}1&\frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}\\ \frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}&1\end{pmatrix}\right)$
Since we have identified $\tilde{\alpha}_{1},\sigma_{\tilde{V}},\sigma_{\tilde{W}},\rho$ from the following equations:
$\rho=\frac{\tilde{\alpha}_{1}\sigma_{\tilde{V}}^{2}+\rho_{U\tilde{V}}}{\big(\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}\big)\sigma_{\tilde{V}}}$
$\sigma^{2}_{\tilde{W}}=\tilde{\alpha}_{1}^{2}\sigma^{2}_{\tilde{V}}+\sigma^{2}_{U}+2\rho_{U\tilde{V}}\tilde{\alpha}_{1}\sigma_{\tilde{V}}\sigma_{U}$
We can identify $\sigma_{U}$ and $\rho_{U\tilde{V}}$, and thus, recover all the structural parameters.
@@ 4. Sharp testable equalities
This section presents the sharp testable equalities for the models described in Section 2. Deviations from these equalities imply violations of the “null hypothesis” that the Tobit model (linear latent index, treatment/instrument independence, instrument relevance, and normality) is valid. The equalities are conditional on parameters being identified under the model’s assumptions, and a discussion on identification is postponed to Section 3.
@@ 4.1. Classic Tobit model
$=$ $P(c_{0}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d\leq U\leq c_{1}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)$
$=$ $\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)-\Phi\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
where $\Phi(.)$ is the standard normal CDF. The first equality follows from the law of total probability. The second through fourth equalities follow from the model’s structure described in Equation 2.4, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y$ for $Y^{*}>0$. The fifth one is given by the latent linear model structure of $Y^{*}$, and finally, the last step follows from assumptions 1 and 2.
Recalling the accumulation point, the observed event of $Y=0$ has a probability,
$P(Y=0|D=d)=P(Y^{*}\leq 0|D=d)=\Phi\left(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right).$ (4.2)
The equalities described in equations 4.1-4.2 fully characterize the distribution of $Y$ conditional on $D$, connecting the probabilities in the observed data to those implied by the Tobit model. We collect these results more formally in Theorem 1.
Theorem 1.
[…]
The proof for the equalities follows from the discussion above, and further details about sharpness are presented in Appendix A. Then, the equalities (4.1)-(4.2) are sharp testable implications for the validity of the classic Tobit Model in 2.4 coupled with assumptions 1-2. They serve as the basis for the test procedure described in Section 5.
Proposition 1 (Non-learnability).
The testable implications and sharpness discussed above show that the classic Tobit model is generally refutable. However, the model is non-verifiable in the sense that we can always construct a joint probability law of $(\bar{Y},D,\bar{U})$ that violates the Tobit model validity but satisfies equalities (4.1)-(4.2). See Appendix E for the proof.
Proposition 2 (Extensions).
The previous derivation can be adjusted for different variations of Tobit models, such as generalizations of the distributional assumptions (barros2018generalized), different thresholds (carson2007Tobit), dynamic Tobit models, or including individual-specific effects (wooldridge2005simple; honore2000estimation; honore1993orthogonality). In particular, note that similar approaches to the ones proposed in Theorem 1 can be used even if the latent errors don’t follow a normal distribution and $Y^{*}$ does not have a linear index form, as long as the model is identified. In appendix C, we derive the equalities for the type 2 Tobit model, and similar logic can be applied to other two-part models. Additional testable results for the aforementioned models are discussed in Appendix D.
@@ 4.2. IV Tobit model
We turn our attention to the IV Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4. The observed data includes $(Y,D,Z)$, and we characterize the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications. The mapping from observed probabilities to their corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$. Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$ (4.3)
- $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
$+$ $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
@@
$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho\right)$ (4.4)
- $\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho\right)$
@@
$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)-\Phi_{W}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z)$ (4.5)
- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
$+$ $\Phi_{W,V}(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
@@
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)-\Phi_{V}(\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.6)
- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
$+$ $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$
@@
$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z)$ (4.7)
- $\Phi_{W,V}(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{0}}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma_{1}}}{\sigma_{\tilde{V}}}z;\rho)$